# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 5 Keystroke Feature Engineering Approaches
Throughout the following section the impact of different keystroke feature engineering techniques is described. The performances of keystroke authentication algorithms depend largely on keystroke data/keystroke features that are used for authentication. So it is important to determine which type of data/feature can give the best result.
@@ 5.1 Effects of Text Filtering
Huang et al. [61] investigated the effect of gibberish text on the performance of keystroke dynamics. Noise or gibberish text can originate when a user plays a computer game or does some unnecessary task. Huang et al. decided that the unnecessary text should be filtered out due to its negative impact on authentication performance. For the evaluation of their hypothesis, they used a novel dataset that was collected when the users were performing normal day-to-day activities. There was no restriction or control over the users. They were free to type anything in order to analyze the noise keystrokes. Huang et al. used a keystroke logger to collect the data of 60 users in a campus student computer labatory. Each user provided at least 11000 keystrokes. They used Gunnetti and Picardi’s algorithm [52] for evaluation. Huang et al. used samples of 1000 keystrokes to create user profile. Each user profile consisted of 10 samples with the rest reserved for testing. They also used the "zone-of-acceptance" algorithm by Leggett et al. [71]. From the dataset, Huang et al. detected four types of gibberish keystrokes - repeating characters, gaming patterns, long string with few distinct characters, long string with no white spaces or separators. They then used two types of algorithms in order to remove gibberish keystrokes. The first one was based on regular expressions which identified four patterns of gibberish keystrokes. The second one was a spell checker which checked for correct English words. They also used context-based filtering and determined that 23.3% data of their dataset was gibberish. The evaluation showed that though gibberish keystrokes did not have much effect on FAR, they negatively impacted FRR. Spell checker filtering showed better performance than regular expressions but the best result was achieved by combining both filters.
@@ 5.2 Effect of Data Size
Huang et al. [59] analyzed the preformance degradation of keystroke authentication systems by experimenting with datasize. In this work, Huang et al. analyzed the impact of reference profile size and test sample. They used two algorithms for their evaluation process. The first one was Gunetti and Picardi’s algorithm [52]. The second one was the "zone-of-acceptance" algorithm by Legget et al. [72]. Huang et al. used this dataset [122] for their evaluation process and performed two experiments. In experiment 1, Huang et al. detected all instances of digraphs and trigraphs in the previously collected user’s data, then created a reference profile and a test sample of specific size. Every instance of data contained within the reference profile was divided into five samples in order to implement the Gunetti-Picardi algorithm. The dataset contained the data of 39 users. Huang et al. performed the tests 50 times to avoid errors. In experiment 2, each user’s data was divided into groups of 1000 digraphs. The range of reference profile size was from 6000 digraph instances to 17000 digraph instances. Only 25 users had enough data to perform the tests. From experiment 1, the results indicated that the FAR value improved when the reference profile size was contant and the test sample size increased. The results also showed that keeping the test sample size constant and increasing the reference profile size did not necessarily improve the FAR. The IPR (Imposter Pass Rate) improved when the reference profile size increased. The IPR values did not decrease for smaller test sample. In experiment 2, Huang et al. used the "zone-of-acceptance" algorithm. The results showed that both FAR and IPR improved with an increasing the reference profile size. Huang et al. also found out that the reference profile should consist of a minimum of 10000 keystrokes or more to achieve good performances.
@@ 5.3 Impact of Faulty Users’ Data
In [89], Ozbek’s goal was to improve the classification performance by removing the keystroke data of the users who degrade the performance. Ozbek used two benchmarking databases with different number of users and passphrases for the evaluation. 1) GREYC Keytroke Data: 133 participants that typed the passphrase "greyc laboratory" on an AZERTY keyboard for the collection. There were 7555 attempts [45]. For this dataset, Ozbek removed some of the users with a low number of captures leaving only the 100 participants with more than 50 attempts. 2)CMU Keystroke Data: 51 participants typed the password ".tie5Roanl" with 400 attempts per user [66]. The data of the two datasets were used as they were and no features were removed. Also no dimension reduction techniques were implemented to improve the performance. To evaluate the preformance Ozbek used support vector machine (SVM), decision tree and K-nearest neighbour (KNN) classifiers. For these three classifiers, each keystroke datum was divided into training, testing and validation set. 70% of the data was used for training, 15% was used for testing and 15% was used for validation. No crossvalidation technique was applied. Ozbek used a histogram to detect the faulty users. Re-classification of the data was performed with the remaining users. For both datasets, Ozbek achieved higher accuracy by eliminating the users with misclassifications at the time of training. First Ozbek removed the worst five users which improved the performance. Ozbek achieved better results by removing more users having misclassifications. However, removing the users from the database was not a standard option. The main goal was to demonstrate that the worst users could be recognized and eliminated if required.
@@ 5.4 Impact of Non-Conventional Keystroke Features
Alsultan, Warwick, and Wei [13] used non-conventional keystroke features for authentication of users. In the case of the non-convention features, the focus was mainly on the overall typing patterns. Percentage of performing certain actions were considered, such as editing actions, general typing actions etc. The goal was to understand the typing behavior of a user. Two main features were considered - semi-timing and editing features. Semi-timing features were different from the common timing features used in previous experiments. These features were calculated for greater amounts of time. The semi-timing features that they used were - 1) Word-per-Minute (WPM) = Number of words / Total typing time in minutes, 2) negUD (negative Up-Down) = Number of negative UDs / Total number of keypairs, 3) negUU (negative UP-Up) = Number of negative UUs / Total number of keypairs. Editing features were characteristics such as typing error frequency, text editing etc. The editing features were - 1) Error rate, 2) CapsLock usage, 3) Shift key usage. Thirty participants provided data for this study and participants had to complete eight typing tasks. The tasks included copying text of 1000 characters. The text was taken from the Guardian newspaper. In the text, there were numbers, both upper and lower case letters and punctuation marks. Alsultan, Warwick, and Wei used a C++ language based GUI program for the data collection. The participants could download the application in their own machines. As the user’s profile, a feature vector of nine features was generated and saved in the database. Out of eight typing tasks each were accompklished with a typing sample from a single source. Features were identified and extracted independently of each typing task.
@@
In the analysis phase, eight samples per subject were used for training and testing of the classifier. As a classifier, decision trees were used in this research. Cross-validation was used in the classification process. In their experiment, Alsultan, Warwick, and Wei used eight samples to perform eight cross-validation experiments. They used seven samples for training and one sample for testing. They used the statistics toolbox in Matlab and two error rates to represent the results - False Accept Rate (FAR) and False Reject Rate (FRR). They obtained low error rates in this study and computed the rates of error for a variable number of participants. For the decision tree classifier, they obtained a FAR of 0.007, 0.0104 and 0.0109 for 15, 25 and 30 participants respectively. Alsultan, Warwick, and Wei obtained a FRR of 0.1, 0.25 and 0.28 for 15, 25 and 30 participants respectively. They also used Support Vector Machines (SVMs) for the classification. Feature subset selection was used in SVMs based classification in order to decrease the dependency between the features.
@@
Alsultan, Warwick, and Wei achieved the best results for a feature set of five features - a FAR of 0.0183 and a FRR of 0.444. They computed the error rates for 30 participants. The focus was on true intruder’s detection in the second part of this study. In this part, typing samples from unknown users were used to test the system. Here, they used binary classification. For the binary classification, they used the training data from 25 legal users and testing data from five intruders. Each intruder generated three typing samples to test the system. For the decision tree classifier, Alsultan, Warwick, and Wei achieved a FAR of 0.011 and a FRR of 0.375 for 25 genuine users without any intruders. For the SVMs, they achieved a FAR of 0.0112 and a FRR of 0.49. When the data of the five intruders was tested against the genuine users’ data, they got the same FAR in case of SVMs. In this study, the FRR improved with a value of 0.28 by utilizing non-conventional features. The lowest FAR was 0.011 which was also comparable to the results generated by conventional features.
@@ 6 Keystroke Authentication on Touch Screen and Mobile Devices
As people are being more dependent on touch screen devices like tablets and mobile phones, it is necessary to implement security and authentication systems based on these devices. This section describes various aspects of keystroke authentication on touch screen and mobile devices.
6.1 Impact of Device Models and Orientations
In [38], El-Abed, Dafer, and Rosenberger proposed a benchmark for keystroke dynamics using a mobile phone and tablet. The touch screen based benchmark was developed in both portrait and landscape orientation. The objective of this benchmark was to evaluate the keystroke authentication algorithms for varying models of devices and different orientations. To observe the acquired keystroke signals, an online visualizer was provided. They used an Android application for developing their benchmark. The app was written in Java for Android devices. They used Nexus 5 and Samsung Galaxy Note 10.1 2014 tablet for data collection. There were two main purposes of this keystroke benchmark. Firstly, they wanted to investigate the patterns of tping of the users for a veriety of orientations, but the device and the password were same. Secondly, they wanted to analyze the effect of different devices on the typing patterns of the users keeping the orientation and the password constant. Data was gathered from 47 users. The users type the password "rhu.university" in four configurations in each session: phone/portrait, tablet/portrait, phone/landscape, tablet/landscape. The key features of the benchmark were - PP, PR, RP, RR, TT (time of typing the password), Screen Orintation (portrait or landscape), Screen Size. El-Abed, Dafer, and Rosenberger performed a statistical analysis using their benchmark and used the Kruskal-Wallis test [55] for the analysis. The goal was to reflect the changes of the keystroke features due to different orientations and different devices. The results indicated that the keystroke features varied significantly for different orientations and different devices. These results must be considered for touch-screen-based mobile authentication technology.
6.2 Impact of Different Types of Features
6.2.1 Combination of Touch-Based and Timing Features
Gautam and Dawadi [42] tried to analyze different features of the touch-screen keyboard for keystroke authentication. They tried to combine the effect of touch-based and timing features and proposed a dataset of 7 users. The authentication system consisted of three steps - Enroll, Verify and Identify. The user authorization process consisted of two phases. For both the training and enrollment phase, a reference template was generated. In the testing or authentication phase, the test samples were matched against the reference template using the Median Vector Classifier [6]. They used four features for the authentication - Key hold time (H), Flight time (FT), Pressure (P), Area (A). They used the string .pie7Crawl for their dataset. The dataset consisted of 47 features. In order to perform their experiment, Gautam and Dawadi utilized a OnePlus 3 Android phone with an AMOLED capacitive touch screen and 1080 * 1920 pixels (5.5 inches). All the 7 participants were touch sceen smart phone users. Every user typed the string .pie7Crawl7 on the touch screen keyboard. Among the 7 participants, one was considered a legal user and the other 6 participants were considered imposters. The legal user participated in 40 sessions. Among those sessions, 10 were reserved for training and the rest for testing. Each of the other 6 participants (imposters) participated in 5 sessions. By using the median vector proximity algorithm, Gautam and Dawadi achieved an average EER of 8.33% and standard deviation EER of 7.07%.
6.2.2 Interaction of Touchscreen Features
In [15], Antal, Szabo, and Laszlo used their dataset of 42 individuals to investigate the interaction of touchscreen features on the performance of keystroke authentication systems. For data collection purpose, they developed an Android application for a Nexus 7 tablet and a Mobil LG Optimus L7 P710. The application had its own software keyboard. During the registration period, the users had to enter their data such as their gender, date of birth, experience of smartphone usage etc. Data was collected in multiple sessions. Most users participated in two sessions in two weeks. The users entered the same password (tie5Roanl) during each session 30 times. A total of 42 individuals provided their data with 51 entries for each user. Entries containing deletions were discarded from the dataset. As all users entered the same password, the data of each user could be used both as legal user and impostor. Among the 42 users, 37 were tablet users and 5 were mobile phone users. The system saved both the timing and touchscreen features. They used WEKA (version 3.6.11) [88] software for their experiment. Weka’s search methods were used to optimize some of the default parameters of the classifiers. They used Bayesian network, Nearest neighbors (k-NN, IBk in Weka), Naive Bayes, Decision trees, Support vector machines, Multilayer perceptrons etc. for the evaluation process. In the identification phase, Antal, Szabo, and Laszlo used two datasets to compute the accuracy - one with touchscreen features (pressure and finger area) and one without touchscreen features. There were 41 features in the first one and 71 features in the second. The data was used without any modification or feature selection and without boosting or tuning methods. They used 10 fold cross-validation in their experiment and computed the classification accuracies by taking the average of 10 runs. All algorithms performed better on the dataset containing touchscreen features. In the verification phase, R script developed by Killourhy and Maxion [66] was used for performing the measurements. Three anomaly-detection algorithms were included in the R script - Euclidean, Manhattan and Mahalanobis metrics. The normalized data was split into three equal segments. In each part, there were 17 samples from each user. User’s template was generated using two-thirds of the data with the remaining data used for FRR testing. In the case of FAR testing, the first five samples from each user were used as impostor data. The measurements were performed three times using a different fold each time for training and testing. The lowest EER was 12.9% which was achieved using the Manhattan metric for both timing and touchscreen features. For timing features, the lowest EER was 15.3%, achieved by the Manhattan metric. Their findings showed that in the identification phase, the accuracy of each classifier increased by 10% due to the addition of touchscreen features. In the verification phase, the EER was reduced by 2.4%.
6.2.3 Two New Features - Speed and Distance
Kambourakis et al. [63] developed and evaluated a touchstroke based authentication system in the Android platform. They used two new features for their experiment apart from commom keystroke features - speed and distance. They adopted two commonly used authentication procedures. In the first procedure, the users had to type a password and in the second case, the users had to type a passphrase. 20 users participated in the data collection between the ages of 19 and 21 years old. All users had a touchscreen based Android smartphone. Kambourakis et al. implemented a prototype of their proposed touchstroke based authentication system in Google’s Android OS. In order to generate the typing behavioral profile of the user and authenticate the user, Kambourakis et al. implemented Waikato Environment for Knowledge Analysis (Weka) as the classification engine. There were two main analysis sub-systems in their mechanism. 1) Enrollment Process: here the typing profile of a user was created and used to train the classifiers, 2) Authentication Process: here the legitimacy of a user was checked. The typing profile of a user was saved in a database. They considered three machine learning algorithms for their experiment - Random Forest, k-NN and MLP. However, MLP was rejected due to memory limitation of smartphone.
For the password scenario, Kambourakis et al. achieved the best results using the Random Forest classifier and first data analysis procedure. For the passphrase scenario, k-NN performed better using the second data analysis procedure. For the password scenario, the average FAR%, FRRand EERvalues were 12.5, 39.4 and 26 respectively. For the passphrase scenario, the average FAR%, FRRand EERwere 23.7, 3.5 and 13.6 respectively. They performed their experiment on a Sony Ericsson Xperia ray. The device had 1 GHz CPU processor, 512 MB RAM, 3.3 inches touchscreen and the Ice Cream Sandwich Android OS. During the classification, 100% of CPU was occupied by the machine learning algorithms. This percentage varied between 91% and 100% at the time of training. The results showed that the first data analysis procedure needed more memory to perform as it generated more data than the second one. Both algorithms were able to perform directly on the device and it took less than one second to authenticate the user.
6.2.4 Novel Sensor Features
Giuffrida et al. [48] proposed a novel keystroke authentication system enhanced with sensor data for authentication of mobile devices. They used novel sensor features to analyze the typing behavior of a user. Machine learning techniques were used to authenticate users and to associate the timing-based keystroke features with movement sensors information. Giuffrida et al. developed UNAGI, a fixed-text authentication system for Android. Their current prototype modified the traditional Android software keyboard and leveraged a number of modules e.g. feature extraction, training and detection. The sensor-enhanced keystroke dynamics techniques were implemented through these support modules. The user had to type a fixed-text password during an authentication session. Their authentication system processed this password for analysis. When the user typed, UNAGI recorded the key-press events and also sampled movement sensor data periodically. They used two Android sensor sampling interfaces: TYPE_LINEAR_ACCELERATION and TYPE_GYROSCOPE. Sensor values were collected at a high sampling frequency. In order to accomplish this, the SENSOR_DELAY_FASTEST flag was specified at sensor listener registration time. The feature extraction module of UNAGI processed the collected data and extracted suitable features. Feaures from the training sessions were fed to the training module to develop a user profile. The detection module compared the features of a testing session against the user profiles for authentication of genuine users. The movement sensor features considered by UNAGI were - root mean square, number of local maxima and minima, minimal and maximal value, mean delta, sum of negative values, sum of positive values, mean value, mean value during KD and KU events, standard deviation. The keytroke features considered by UNAGI were - KD-KD time, KD-KU time, KU-KD time, KU-KU time. After the feature extraction phase, the output was a vector which consisted of all the features - both keystroke and movement sensor features. In their experiment, threshold-based binary classification algorithms were suitable. Giuffrida et al. used one-class SVM, Naive Bayes, k-Nearest Neighbours (kNN), the "mean algorithm". They also used several distance metrics - Euclidean, Euclidean normed, Manhattan, Manhattan scaled and Mahalanobis. They used three different configurations for evaluation purpose: keystroke data only, sensor data only, and combination of both. Results showed better accuracy with sensor-based features rather than keystroke timing features- i.e., an EER of 0.08% for sensor-based features vs. an EER of 4.97% for keystroke timing features.
6.2.5 Touch-Based Features with Three Distinct Hand Postures
In [22], Buschek, De Luca, and Alt proposed a experiment investigating mobile keystroke biometrics. They compared touch-based features in consideration with three distinct hand postures and tried to evaluate touch-specific features for user authentication. They used two types of typing features.
  1. 1.
    Temporal features: hole time, flight time, up-up times, down-down times etc.
  2. 2.
    Spatial touch-specific features: exact touch locations, offsets, touch jumps, drag, touch area sizes, ellipses axes, touch pressure etc.
They used two types of models - 1) anomaly detectors, 2) classification methods. For this study, data was collected in two sessions with a one week gap between sessions. A total of 28 users participated in the data collection and the average age of the users was 25.
A Nexus 5 phone was utilized for the data collection. A unique software based keyboard was used to extract the features. Each user participated in two sessions. Each session included data collected using three seperate hand orientations. The participants were instructed to type 6 distinct passwords 20 times using each hand posture. From their analysis it was observed that models achieved better results on data of a single session - the EERs were less than half of the values obtained for data across sessions. In the case of evaluation, classifiers performed better than anomaly detectors (11.7-31.1% lower EERs). EERs increased by 86.3% for different hand postures. Regarding the features, the best spatial feature sets performed better than the best temporal sets. EERs were reduced by 14.3 - 23.5% using the spatial touch features relative to the traditional temporal features. Also EERs were reduced by using spatial touch features instead of pressure features. The combination of spatial and temporal features provided the best results - EERs were reduced by 8.5 - 26.3% relative to the best spatial features, and by 26.4 - 36.8% relative to the best temporal features.
@@ 6.2.5 Touch-Based Features with Three Distinct Hand Postures
Table 4: A summary of keystroke dynamics on touch screen and mobile devices. In this table, algorithms used for keystroke authentication on touch screen and mobile devices are mentioned along with the data on which the algorithms were applied. Also the results or contributions of the reference papers are provided.

### v3

@@ 5 Keystroke Feature Engineering Approaches
Throughout the following section, the impact of different keystroke feature engineering techniques is described. The performances of keystroke authentication algorithms depend largely on keystroke data/keystroke features that are used for authentication, so it is important to determine which features give the best results.
@@ 5.1 Effects of Text Filtering
Huang et al. [61] investigated the effect of gibberish text on the performance of keystroke dynamics. Noise or gibberish text can originate when a user plays a computer game or does some unnecessary task. Huang et al. decided that the unnecessary text should be filtered out due to its negative impact on authentication performance. For the evaluation of their hypothesis, they used a novel dataset that was collected when the users were performing normal day-to-day activities. There was no restriction or control over the users. They were free to type anything in order to analyze the noise keystrokes. Huang et al. used a keystroke logger to collect the data of 60 users in a campus student computer laboratory. Each user provided at least 11000 keystrokes. They used Gunnetti and Picardi’s algorithm [52] for evaluation. They used samples of 1000 keystrokes to create a user profile. Each user profile consisted of 10 samples with the rest reserved for testing. They also used the “zone-of-acceptance” algorithm by Leggett et al. [71]. From the dataset, they detected four types of gibberish keystrokes: repeating characters, gaming patterns, long strings with few distinct characters, and long strings with no white spaces or separators. They then used two types of algorithms in order to remove gibberish keystrokes. The first one was based on regular expressions that identified four patterns of gibberish keystrokes. The second one was a spell checker that checked for correct English words. They also used context-based filtering and determined that 23.3% data in their dataset was gibberish. The evaluation showed that though gibberish keystrokes did not have much effect on FAR, they negatively impacted FRR. Spell checker filtering showed better performance than regular expressions, but the best result was achieved by combining both filters.
@@ 5.2 Effect of Data Size
Huang et al. [59] analyzed the performance degradation of keystroke authentication systems by experimenting with data size. They analyzed the impact of reference profile size and test sample. They used two algorithms for their evaluation process. The first one was Gunetti and Picardi’s algorithm [52]. The second one was the “zone-of-acceptance” algorithm by Legget et al. [72]. They used the dataset developed by Vural et al. [122] for their evaluation and performed two experiments. In the first experiment, they detected all instances of digraphs and trigraphs in the previously collected user’s data, then created a reference profile and a test sample of a specific size. Every instance of data contained within the reference profile was divided into five samples in order to implement the Gunetti-Picardi algorithm. The dataset contained the data of 39 users. They performed the tests 50 times to avoid errors. In the second experiment, each user’s data was divided into groups of 1000 digraphs. The range of reference profile size was from 6000 digraph instances to 17000 digraph instances. Only 25 users had enough data to perform the tests. From the first experiment, the results indicated that the FAR value improved when the reference profile size was constant and the test sample size increased. The results also showed that keeping the test sample size constant and increasing the reference profile size did not necessarily improve the FAR. The IPR (Imposter Pass Rate) improved when the reference profile size increased. The IPR values did not decrease for smaller test samples. In the second experiment, Huang et al. used the “zone-of-acceptance” algorithm. The results showed that both FAR and IPR improved with an increasing the reference profile size. They also found out that the reference profile should consist of a minimum of 10,000 keystrokes or more to achieve good performances.
@@ 5.3 Impact of Faulty Users’ Data
Ozbek’s goal was to improve the classification performance by removing the keystroke data of the users who degrade the performance [89]. Ozbek used two benchmarking databases with different numbers of users and passphrases for the evaluation.
  1. 1.
    GREYC Keystroke Data: 133 participants typed the passphrase “greyc laboratory” on an AZERTY keyboard for the collection. There were 7555 attempts [45]. For this dataset, Ozbek removed some of the users with a low number of captures leaving only the 100 participants with more than 50 attempts.
  2. 2.
    CMU Keystroke Data: 51 participants typed the password “.tie5Roanl” with 400 attempts per user [66].
The data of the two datasets were used as they were and no features were removed. Also, no dimension-reduction techniques were implemented to improve the performance. To evaluate the performance Ozbek used support vector machine (SVM), decision tree, and K-nearest neighbor (KNN) classifiers. For these three classifiers, each keystroke datum was divided into training, testing, and validation sets. 70% of the data was used for training, 15% was used for testing and 15% was used for validation. No cross-validation technique was applied. Ozbek used a histogram to detect the faulty users. Re-classification of the data was performed with the remaining users. For both datasets, Ozbek achieved higher accuracy by eliminating the users with misclassifications at the time of training. First Ozbek removed the worst five users which improved the performance. Ozbek achieved better results by removing more users having misclassifications. However, removing the users from the database was not a standard option. The main goal was to demonstrate that the worst users could be recognized and eliminated if required.
@@ 5.4 Impact of Non-Conventional Keystroke Features
Alsultan, Warwick, and Wei [13] used non-conventional keystroke features for the authentication of users. In the case of the non-convention features, the focus was mainly on the overall typing patterns. The percentage of performing certain actions was considered, such as editing actions, general typing actions, etc. The goal was to understand the typing behavior of a user. Two main features were considered - semi-timing and editing features. Semi-timing features were different from the common timing features used in previous experiments. These features were calculated for greater amounts of time. The semi-timing features that they used were: 1) Word-per-Minute (WPM) = Number of words / Total typing time in minutes, 2) negUD (negative Up-Down) = Number of negative UDs / Total number of keypairs, 3) negUU (negative UP-Up) = Number of negative UUs / Total number of keypairs. Editing features were characteristics such as typing error frequency, text editing, etc. The editing features were: 1) Error rate, 2) CapsLock usage, and 3) Shift key usage. Thirty participants provided data for this study and participants had to complete eight typing tasks. The tasks included copying the text of 1000 characters. The text was taken from the Guardian newspaper. In the text, there were numbers, both upper and lower case letters, and punctuation marks. They used a C++ based GUI program for the data collection. The participants could download the application on their own machines. As the user’s profile, a feature vector of nine features was generated and saved in the database. Out of eight typing tasks, each was accomplished with a typing sample from a single source. Features were identified and extracted independently of each typing task.
@@
In the analysis phase, eight samples per subject were used for training and testing of the classifier. As a classifier, decision trees were used in this research. Cross-validation was used in the classification process. In their experiment, Alsultan, Warwick, and Wei used eight samples to perform eight cross-validation experiments. They used seven samples for training and one sample for testing. They used the statistics toolbox in Matlab and two error rates to represent the results - False Accept Rate (FAR) and False Reject Rate (FRR). They obtained low error rates in this study and computed the rates of error for a variable number of participants. For the decision tree classifier, they obtained a FAR of 0.007, 0.0104, and 0.0109 for 15, 25, and 30 participants respectively. They obtained a FRR of 0.1, 0.25, and 0.28 for 15, 25, and 30 participants respectively.
@@
The focus was on true intruder detection in the second part of this study. In this part, typing samples from unknown users were used to test the system. Here, they used binary classification. For the binary classification, they used the training data from 25 legal users and the testing data from five intruders. Each intruder generated three typing samples to test the system. For the decision tree classifier, they achieved a FAR of 0.011 and a FRR of 0.375 for 25 genuine users without any intruders. For the Support Vector Machines (SVMs), they achieved a FAR of 0.0112 and a FRR of 0.49. When the data of the five intruders was tested against the genuine users’ data, they got the same FAR in the case of SVMs. In this study, the FRR improved to a value of 0.28 utilizing non-conventional characteristics. The lowest FAR was 0.011, which was also comparable to the results generated by conventional features.
@@ 6 Keystroke Authentication on Touch Screen and Mobile Devices
As people are increasingly dependent on touch screen devices such as tablets and mobile phones, it is necessary to implement security and authentication systems based on these devices. This section describes various aspects of keystroke authentication on touch screen and mobile devices; a summary is shown in Table 4.
@@
Table 4: A summary of keystroke dynamics on touch screen and mobile devices. In this table, algorithms used for keystroke authentication on touch screen and mobile devices are mentioned along with the data on which the algorithms were applied. Also, the results or contributions of the reference papers are provided.